From GlueXWiki
Jump to: navigation, search

April 8, 2010 FDC meeting

Tentative Agenda

  1. Off-site production space
  2. Electronics (Fernando)
  3. Cathode redesign redesign, central hole (Bill, Roger)
  4. Redesigned frame prototype
    • plans for tests
  5. Full scale prototype (Beni)
    • two-layer HV tests
    • third layer status
    • preparation for cosmics tests
  6. Other


Participants: Eugene, Bill, Roger, Brian, Simon, Glenn, Casey, Beni, Fernando, Lubomir

Off-site production space

- Eugene informed us that the property at Blue Crab will be rented soon. JLab will pay for the rent and common infrastructure. We will be allowed to overspend to pay for our infrastructure (clean space, ...).

- Bill presented the floor plan for our space: 2000 sq.ft clean room, plus ~1000 sq.ft semi-clean and other areas (see the above link). Brian suggested replacing areas 2 and 5 (soldering and assembling/testing). Eugene proposed to separate the wire stringing (if at JLab) from the other soldering places. The cleaning area was discussed: using alcohol, DI water, or ultrasonic. Fernando suggested to make a diagram of the production flow. Glenn informed us about the expected timing (most likely in October we can start using the space) and that the cost of the clean area would be $70/sq.ft. We have to specify the power consumption (maximal, duty cycle, power dissipated inside the clean room), maximum number of people working there, description of their work. Lubomir is responsible for that information.


- next week the final PCB design will be posted and everybody is encourage to look at it

- we discussed if spare cables are needed and whether spares have to be installed together with the used cables. People have different experience at different experiments. We decided to include 5% spare cables (as "extra process equipment") in the procurement, while all the other issues can be discussed later.

Cathode redesign

- Roger has implemented most of the changes shown in the above link. Some additional small modifications were discussed. Beni's point about the central hole was accepted: rounded strips were OK in the full scale prototype and we don't want to change this.

Redesigned frame prototype

- the newly designed frames are ready; we started testing the cathode sandwich. Bill explained what he wants from these tests: to find the smallest flow rate at which we can bubble the chamber (i.e. to have ~30 Pa inside). Then, to repeat the same with the redesigned wire frame also mounted. We discussed whether it is needed to include the prototype as a fourth layer during the cosmics tests. It will require some additional work.

Full scale prototype

- the two layers are assembled together and operational, in the sense that can hold the nominal HV of 2200V with 40/60 Ar/CO2. Beni explained that after we cleaned the first layer twice and the second layer once, they performe much better.

- Casey was working on fixing the third wire layer and reported about the wire problems there: all the wires are not touching the soldering pads, some wires not fully covered with solder, most of the wire tips not cut properly, damaged connection between the epoxy and the soldering pad (normally there was a previous attempt to fix these places), three wires for now can't be repaired without replacing the whole wire. We decided to fix what's easy and to put the third layer "as it is" for the tests with cosmics.

- The plan is to start testing the three layers next week with cosmics in horizontal position. This would require some additional mechanical support. Beni started already working with the scintillators and the IU chamber in the testing set-up. He recently got the CAEN HV frame back from repair. Apart from one fan and a dent on the frame, it's working fine.